OpenStudy (anonymous):

Name FIVE x-values for which the tangent function equals 0.

OpenStudy (anonymous):

find nine values of \(x\) for which \(\sin(x)=0\) here is my example : \(\pi\) since \(\sin(\pi)=0\)

OpenStudy (anonymous):

so like sin(o), sin(180) , sin(-180)

OpenStudy (anonymous):

@satellite73